Please help me to obtain the derivative. for given: Y=x^2-4X+5
I think the original points are: (X1, X2-4x+5)
and second point should be (x + h, (X + h)^2
next I need to find oy and ox (which = the y2 + y1 and x1 + x2)
This is where I get confused, I don't know what to do next.

Answer by stanbon(75887)

derivative. for given: Y=x^2-4X+5
-------------
limit as h goes to zero of [f(x+h)-f(x)]/[h]
-------------
f(x+h) = (x+h)^2-4(x+h)+5
= x^2+2hx+h^2-4x-4h+5
-----------
f(x+h)-f(x) = 2hx+h^2-4h
-----------
[f(x+h)-f(x)]/[h] = 2x+h-4
--------------
limit as h goes to zero is 2x-4 (this is the derivative you want)
==================
Cheers,
Stan H.